OpenStudy (anonymous):

If the arc on a particular circle has a length of 12 units, and the circumference of the circle is 36, which is the degree measure of the arc?

OpenStudy (anteater):

First off, what is the ratio of the length of the arc to the total circumference?

OpenStudy (jh3power):

12/36=x/360 x=120

OpenStudy (anteater):

Yes :D

OpenStudy (anonymous):

thanks

OpenStudy (anteater):

Wlcm!!!

OpenStudy (jh3power):

welcome